Please help with this exercise, greatly thanks if you could do in in details so that i can understand how the numbers come from.

The investor Comedy acquires a 30% interest in mother Co. Drama for $800,000

Assume that the investor has the ability to significantly influence the investee

2014: Drama report net losses of $240,000 and issued bonds for $120,000

2015: Drama report net income of $600,000 and pays dividend for $300,000

2016: Drama reports net losses of $120, 000. At the end of the year, Comedy sells one fourth of its investment in Drama for $200,000

Determine of the value of the investment at the end of each year
Answer #1

Determination of Value of Investment

Total value of Drama = $800000 / 30 % = $ 26,66,667

Value of investment at the end of each year

Year 2014

Value of investment = Value of investment at the beginning of year - Value of bonds issue - Net loss

= $26,66,667 - $120,000 - $240,000

= $ 23,06,667

Value of investment of Comedy in Drama = $23,06,667 * 30% = $692,000

Year 2015

Value of investment = $23,06,667 + $600,000 - $ 300,000

= $ 26,06,667

Value of investment of Comedy in Drama = $26,06,667 * 30% = $692,000 = $ 782,000

Year 2016

Value of investment = $ 26,06,667 - $ 120,000 - $13,500 ( loss on buy back of investment )

= $ 24,73,167

Value of investment of Comedy in Drama = $ 24,73,167 * 30% * 0.75  = $556,462.58